# Jackknife Empirical Likelihood Method for U Statistics Based on Multivariate Samples and its Applications

### v1

@@ 2 JEL for three multivariate sample
Therefore, $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})=(\mathbb{X}_{1},\dots,\mathbb{X}_{n_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{n_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{n_{3}})$, where $\mathbb{W}_{i}^{\prime}s$ are independent but need not be identically distributed. Then, the U-statistics based on single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ with degree $m=m_{1}+m_{2}+m_{3}$ is given as follows:
$U_{n}=\frac{1}{{n\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right),$ (9)
@@
=$=\begin{cases}&\text{if}\;\;1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ c^{*}\;h\left(\mathbb{W}_{i_{1}},\dots,\mathbb{W}_{i_{m_{1}}},\mathbb{W}_{j_{1}},\ldots,\mathbb{W}_{j_{m_{2}}},\mathbb{W}_{k_{1}},\dots,\mathbb{W}_{k_{m_{3}}}\right),&n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ &n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\\ 0,&\text{otherwise}\end{cases}$
@@
Consider the U-statistics after deleting observation $\mathbb{W}_{l}$, from sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$, given by
$U_{n-1}^{(-l)}=\frac{1}{{n-1\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\\ l_{i}\neq l,\;i=1,2,\dots,m\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right),\;\;\forall\;l=1,2,\dots,n.$ (10)
Define the jackknife pseudo-values $V_{l}$, for combined single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ as
$V_{l}=nU_{n}-(n-1)U_{n-1}^{(-l)},\;\;\forall\;l=1,2,\dots,n.$ (11)
It is easy to verify that $U_{n}=\frac{1}{n}\sum_{i=1}^{n}V_{i},$ where $V_{i}$ ’s are asymptotically independent random variables ([21]).
The expected values of $V_{l}^{\prime}s$ are given below (provided in Appendix A).
$EV_{l}=\begin{cases}\theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right],&\text{ if }\;l=1,2,\ldots,n_{1}\\ ~{}\\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right],&\text{ if }\;l=n_{1}+1,n_{1}+2,\ldots,n_{1}+n_{2}\\ ~{}\\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right],&\text{ if }\;l=n_{1}+n_{2}+1,\ldots,n\end{cases}.$ (12)
@@
$L(\theta)=\max\left\{\prod_{i=1}^{n}p_{i};\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$ (13)
@@
$R(\theta)=\frac{L(\theta)}{n^{-n}}=\max\left\{\prod_{i=1}^{n}\left(np_{i}\right);\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$ (14)
@@
$p_{i}=\frac{1}{n}\frac{1}{1+\lambda\left(V_{i}-EV_{i}\right)},\;\;\forall\;\;i=1,2,\dots,n,$ (15)
@@
$\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}=0.$ (16)
@@
We can use $\log R(\theta)$ for constructing confidence intervals or developing test on $\theta$. For this we need to obtain the asymptotic distribution of $logR(\theta)$. We next derive the limiting distribution of the jackknife empirical log-likelihood ratio as an analogous to Wilk’s theorem. The proof of the following theorem is given in the Appendix B.
@@
Using Theorem 2.2, we construct a $(1-\alpha)$ level confidence interval for $\theta$, given by
$CI=\{\theta\;|\;-2\log R(\theta)\leq\chi^{2}_{1,1-\alpha}\},$
where $\chi^{2}_{1,1-\alpha}$ is the $(1-\alpha)$-th percentile point of a $\chi^{2}$ distribution with one degree of freedom.
2.1 Application: Volumes under ROC surfaces
This subsection presents an application of the results obtained above to the Volume Under the ROC Surface (VUS), introduced by [14]. Traditionally confined to binary classifications, ROC curves and their associated Area Under the Curve (AUC) metrics provide critical insights into diagnostic accuracy but are limited in their applicability to multi-class scenarios. The VUS extends AUC by accommodating three-class diagnostic scenarios. This evolution addresses the inherent limitations of binary classifications in medical diagnostics, where conditions often manifest along a spectrum rather than as discrete categories [24]. [25], [4], and [1] considered non-parametric inference for VUS using the JEL method.
Consider levels of three biomarkers used to assess the risk of developing cardiovascular disease: let $X$ be the level of Biomarker 1 (e.g., LDL cholesterol), $Y$ be the level of Biomarker 2 (e.g., HDL cholesterol), and $Z$ be the level of Biomarker 3 (e.g., triglycerides). We are interested in finding the probability that a patient has an LDL cholesterol level $X$ that is lower than their HDL cholesterol level $Y$, and their HDL cholesterol level $Y$ that is lower than their triglyceride level $Z$, i.e., $P(X<Y<Z)$. This probability, which is the VUS, can be interpreted as a measure of how often the biomarker levels are ordered in a specific way that might be associated with different risk levels or health conditions ([5]).
@@ 2.1 Application: Volumes under ROC surfaces
$\theta=P\left(X_{11}<Y_{11}<Z_{11}\right)-P\left(X_{12}<Y_{12}<Z_{12}\right).$
In the following part of this subsection, we use three different non-parametric methods and construct confidence intervals for $\theta$.
U-Statistics based Confidence Intervals :
We can rewrite $\theta$ as
$\theta$ $=E\left[I\left(X_{11}<Y_{11}<Z_{11}\right)\right]-E\left[I\left(X_{12}<Y_{12}<Z_{12}\right)\right]$
$=E\left[I\left(X_{11}<Y_{11}<Z_{11}\right)-I\left(X_{12}<Y_{12}<Z_{12}\right)\right],$
@@
An estimator of $\theta$ can be found as U-statistic with kernel $h$ as
@@
The U-statistic $U_{n_{1},n_{2},n_{3}}^{VUS}$ is an unbiased and consistent estimator of $\theta$ (Lehmann, (1951)). Also, using Theorem 2.1, a two sided $(1-\alpha)$ level CI for $\theta$ based on the asymptotic distribution of $U_{n_{1},n_{2},n_{3}}^{VUS}$ is
$\left(U_{n_{1},n_{2},n_{3}}^{VUS}-z_{\alpha/2}\hat{\sigma},U_{n_{1},n_{2},n_{3}}^{VUS}+z_{\alpha/2}\hat{\sigma}\right),$
where $\hat{\sigma}$ is the jackknife pseudo-values based consistent estimator of $Var(U_{n_{1},n_{2},n_{3}}^{VUS})$ (see [2]), as defined by (2).
JEL based Confidence Interval :
Now, we use the Theorem 2.2 to construct JEL based confidence intervals for $\theta$. Here with respect to Theorem 2.2, we have $q=2$ and $m_{1}=m_{2}=m_{2}=1$, and on the same lines as of Section 2, we can define $\log R(\theta)$. Therefore, using Theorem 2.2, the JEL based confidence interval for $\theta$ at level $(1-\alpha)$ is given by
$CI=\{\theta\;|\;-2\log R(\theta)\leq\chi^{2}_{1,1-\alpha}\}.$
Kernel Based Confidence Interval :
To estimate the difference of two VUS’s smoothly [8] proposed the kernel-smoothed version of the difference of two VUS’s as follows
$\hat{\theta}=\frac{1}{n_{1}n_{2}n_{3}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\sum_{k=1}^{n_{3}}\left(\Phi\left(\frac{Y_{1j}-X_{1i}}{\sqrt{h_{1}^{2}+h_{2}^{2}}}\right)\Phi\left(\frac{Z_{1k}-Y_{1j}}{\sqrt{h_{2}^{2}+h_{3}^{2}}}\right)-\Phi\left(\frac{Y_{2j}-X_{2i}}{\sqrt{h_{4}^{2}+h_{5}^{2}}}\right)\Phi\left(\frac{Z_{2k}-Y_{2j}}{\sqrt{h_{5}^{2}+h_{6}^{2}}}\right)\right),$
@@
$h_{i}=0.9\min\left\{sd\left(x_{i}\right)\right.,\left.\operatorname{iqr}\left(x_{i}\right)/1.34\right\}n_{i}^{-0.2},\;\;\;\;i=1,2,\dots,6,$
where $sd(\cdot)$ and $\operatorname{iqr}(\cdot)$ represent the standard deviation and the interquartile range. [22] introduced the bandwidth $h_{i}$, which regulates the degree of smoothing. This asymptotic bandwidth selection is effective for a broad range of density functions. Such bandwidths have been utilized by many researchers in diagnostic studies.
To construct the kernel based confidence interval for $\theta$, we employ the non-parametric bootstrap method to obtain the confidence interval. In each bootstrap replication, we find a point estimate $\hat{\theta}$ of $\theta$, that allows us to construct the confidence intervals using the percentile method.
3 Simulation

### v2

@@ 2.2. Three sample multivariate U-Statistics - JEL
Therefore, $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})=(\mathbb{X}_{1},\dots,\mathbb{X}_{n_{1}},\mathbb{Y}_{1},\dots,\mathbb{Y}_{n_{2}},\mathbb{Z}_{1},\dots,\mathbb{Z}_{n_{3}})$, where $\mathbb{W}_{i}^{\prime}s$ are independent but need not be identically distributed. Then, the U-statistic based on single sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ with degree $m=m_{1}+m_{2}+m_{3}$ is given as follows:
$U_{n}=\frac{1}{{n\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right),$ (2.9)
@@
$\begin{cases}&\text{if}\;\;1\leq i_{1}<\ldots<i_{m_{1}}\leq n_{1}\\ c^{*}\;h\left(\mathbb{W}_{i_{1}},\dots,\mathbb{W}_{i_{m_{1}}},\mathbb{W}_{j_{1}},\ldots,\mathbb{W}_{j_{m_{2}}},\mathbb{W}_{k_{1}},\dots,\mathbb{W}_{k_{m_{3}}}\right),&n_{1}+1\leq j_{1}<\ldots<j_{m_{2}}\leq n_{1}+n_{2}\\ &n_{1}+n_{2}+1\leq k_{1}<\ldots<k_{m_{3}}\leq n\\ 0,&\text{otherwise}\end{cases}$
@@
Consider the U-statistic after deleting observation $\mathbb{W}_{l}$, from sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$, given by
$U_{n-1}^{(-l)}=\frac{1}{{n-1\choose m}}\mathop{\sum}_{\begin{subarray}{c}1\leq l_{1}<\ldots<l_{m}\leq n\\ l_{i}\neq l,\;i=1,2,\dots,m\end{subarray}}\tilde{h}\left(\mathbb{W}_{l_{1}},\mathbb{W}_{l_{2}},\ldots,\mathbb{W}_{l_{m}}\right),\;\;\forall\;l=1,2,\dots,n.$ (2.10)
Define the jackknife pseudo-values $V_{l}$, for combined sample $(\mathbb{W}_{1},\mathbb{W}_{2},\dots,\mathbb{W}_{n})$ as
$V_{l}=nU_{n}-(n-1)U_{n-1}^{(-l)},\;\;\forall\;l=1,2,\dots,n.$ (2.11)
It is easy to verify that $U_{n}=\frac{1}{n}\sum_{i=1}^{n}V_{i},$ where $V_{i}$ ’s are asymptotically independent random variables (shi1984approximate).
The expected values of $V_{l}^{\prime}s$ are given below (details are provided in Appendix A).
$EV_{l}=\begin{cases}\theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{1}-m_{1})}{n_{1}}\right],&\text{ if }\;l=1,2,\ldots,n_{1}\\ \penalty 10000\ \\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{2}-m_{2})}{n_{2}}\right],&\text{ if }\;l=n_{1}+1,n_{1}+2,\ldots,n_{1}+n_{2}\\ \penalty 10000\ \\ \theta\left(\frac{n}{n-m}\right)\left[n-m-\frac{(n-1)(n_{3}-m_{3})}{n_{3}}\right],&\text{ if }\;l=n_{1}+n_{2}+1,\ldots,n\end{cases}.$ (2.12)
@@
$L(\theta)=\max\left\{\prod_{i=1}^{n}p_{i};\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$ (2.13)
@@
$R(\theta)=\frac{L(\theta)}{n^{-n}}=\max\left\{\prod_{i=1}^{n}\left(np_{i}\right);\sum_{i=1}^{n}p_{i}=1,\sum_{i=1}^{n}p_{i}\left(V_{i}-EV_{i}\right)=0\right\}.$ (2.14)
@@
$p_{i}=\frac{1}{n}\frac{1}{1+\lambda\left(V_{i}-EV_{i}\right)},\;\;\forall\;\;i=1,2,\dots,n,$ (2.15)
@@
$\frac{1}{n}\sum_{i=1}^{n}\frac{V_{i}-EV_{i}}{1+\lambda\left(V_{i}-EV_{i}\right)}=0.$ (2.16)
@@
We can use $\log R(\theta)$ for constructing confidence intervals or developing test on $\theta$. For this, we need to obtain the asymptotic distribution of $\log R(\theta)$. We next derive the limiting distribution of the jackknife empirical log-likelihood ratio as an analog to Wilk’s theorem. The proof of the following theorem is given in Appendix B.
@@
3. Application: Volumes under ROC surfaces
The receiver operating characteristic (ROC) curve and its associated area under the curve (AUC) are fundamental tools for assessing the diagnostic accuracy of binary classifiers, summarizing their ability to discriminate between two outcome states across all possible threshold values. In medical science, ROC analysis is routinely used to evaluate diagnostic tests, compare biomarkers, and determine optimal thresholds for disease detection (yang2024transformed).
However, in medical diagnostics, disease severity may exist on a continuum—such as healthy, intermediate, and diseased — rather than fitting neatly into a binary framework. For example, consider levels of three biomarkers used to assess the risk of developing cardiovascular disease: let $X$ be the level of Biomarker 1 (e.g., LDL cholesterol), $Y$ be the level of Biomarker 2 (e.g., HDL cholesterol), and $Z$ be the level of Biomarker 3 (e.g., triglycerides). We are interested in finding the probability that a patient has an LDL cholesterol level $X$ that is lower than their HDL cholesterol level $Y$, and their HDL cholesterol level $Y$ that is lower than their triglyceride level $Z$, i.e., $P(X<Y<Z)$. This probability, is called Volume under ROC surface and denoted by VUS. It can be interpreted as a measure of how often the biomarker levels are ordered in a specific way that might be associated with different risk levels or health conditions (he2008).
This extension takes care of the inherent limitations of binary classifications in medical diagnostics, where conditions often manifest along a spectrum rather than as discrete categories. tian2011exact. wang2010empirical, guangming2013nonparametric, and an2018jackknife considered non-parametric inference for VUS using the JEL method.
@@ 3. Application: Volumes under ROC surfaces
$\theta_{V}=P\left(X_{11}<Y_{11}<Z_{11}\right)-P\left(X_{12}<Y_{12}<Z_{12}\right).$
In the following part of this subsection, we use three different non-parametric methods for finding confidence intervals for $\theta_{V}$. We adopt the notations introduced in Sections 2.1 and 2.2, with $q=2$ and $m_{1}=m_{2}=m_{3}=1$.
U-statistic based confidence intervals for $\theta_{V}$:
We can rewrite $\theta_{V}$ as
$\theta_{V}$ $=E\left[I\left(X_{11}<Y_{11}<Z_{11}\right)-I\left(X_{12}<Y_{12}<Z_{12}\right)\right],$
@@
An estimator of $\theta_{V}$ can be found as U-statistic with kernel $h$ as
@@
The U-statistic $U_{n_{1},n_{2},n_{3}}^{VUS}$ is an unbiased and consistent estimator of $\theta$ (Lehmann, (1951)). Also, using Theorem 2.1, a two sided asymptotic $(1-\alpha)$ level CI for $\theta_{V}$ based on the asymptotic distribution of $U_{n_{1},n_{2},n_{3}}^{VUS}$ is
[…]
where $\hat{\sigma}$ is the jackknife pseudo-values based consistent estimator of $Var(U_{n_{1},n_{2},n_{3}}^{VUS})$ (see arvesen1969jackknifing), as defined by (2.1). Note that this confidence interval was also considered in an2018jackknife.
JEL based confidence interval for $\theta_{V}$:
We use Theorem 2.3 to find JEL based confidence intervals for $\theta_{V}$. We can define $\log R(\theta_{V})$ on the same lines as of Section 2.2. Hence, the JEL based confidence interval for $\theta_{V}$ at level $(1-\alpha)$ is given by
$CI=\{\theta_{V}\;|\;-2\log R(\theta_{V})\leq\chi^{2}_{1,1-\alpha}\}.$ (3.2)
Note that JEL based confidence interval was also considered in an2018jackknife.
Kernel based confidence interval for $\theta_{V}$: To estimate the difference of two VUS’s kang2013 proposed the kernel-smoothed version of the difference of two VUS’s as follows;
${\hat{\theta}_{V}}=\frac{1}{n_{1}n_{2}n_{3}}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\sum_{k=1}^{n_{3}}\left(\Phi\left(\frac{Y_{1j}-X_{1i}}{\sqrt{b_{1}^{2}+b_{2}^{2}}}\right)\Phi\left(\frac{Z_{1k}-Y_{1j}}{\sqrt{b_{2}^{2}+b_{3}^{2}}}\right)-\Phi\left(\frac{Y_{2j}-X_{2i}}{\sqrt{b_{4}^{2}+b_{5}^{2}}}\right)\Phi\left(\frac{Z_{2k}-Y_{2j}}{\sqrt{b_{5}^{2}+b_{6}^{2}}}\right)\right),$ (3.3)
@@
$b_{i}=0.9\min\left\{sd\left(x_{i}\right)\right.,\left.\operatorname{iqr}\left(x_{i}\right)/1.34\right\}n_{i}^{-0.2},\;\;\;\;i=1,2,\dots,6,$
where $sd(\cdot)$ and $\operatorname{iqr}(\cdot)$ represent the standard deviation and the interquartile range. silverman2018 introduced the bandwidth $b_{i}$ which controls the level of smoothing in kernel estimation. This asymptotic bandwidth selection method has proven effective across a wide range of density functions and has been widely adopted by researchers in diagnostic studies.
We use the estimator (3.3) and bootstrap method to obtain a kernel-based confidence interval for $\theta_{V}$. In each bootstrap replication, we compute a point estimate $\hat{\theta}_{V}$ of $\theta_{V}$, which allows us to construct the confidence interval using the percentile method. The kernel-based bootstrap confidence interval for $\theta_{V}$ is obtained in Step 5 of Algorithm 1. This approach is commonly used for constructing confidence intervals; see polansky2001bandwidth, feinberg2024kernel, and lin2025bootstrap.
Algorithm 1 Bootstrap Percentile CI for a Kernel-Based Estimator of $\theta_{V}$
1:Compute the original estimate $\hat{\theta}_{V}$ (as defined by (3.3)).
2:for $b=1,\dots,B$ do
3:a: Draw bootstrap samples $(X_{1}^{*(b)},\dots,X_{n_{1}}^{*(b)})$, $(Y_{1}^{*(b)},\dots,Y_{n_{2}}^{*(b)})$ and $(Z_{1}^{*(b)},\dots,Z_{n_{3}}^{*(b)})$ by sampling with replacement from $\{X_{1},\dots,X_{n_{1}}\}$, $\{Y_{1},\dots,Y_{n_{2}}\}$ and $\{Z_{1},\dots,Z_{n_{3}}\}$, respectively.
4:b: Based on these bootstrap samples, compute the bootstrap estimate $\hat{\theta}_{V}^{*(b)}$ using (3.3).
5:Sort $\hat{\theta}_{V}^{*(1)},\dots,\hat{\theta}_{V}^{*(B)}$ in ascending order to obtain
$\hat{\theta}^{*(1)}_{V}\leq\cdots\leq\hat{\theta}^{*(B)}_{V}$.
6:Let
$\hat{\theta}^{*}_{\alpha/2}=\hat{\theta}^{*(\lceil B(\alpha/2)\rceil)}_{V},\quad\hat{\theta}^{*}_{1-\alpha/2}=\hat{\theta}^{*(\lceil B(1-\alpha/2)\rceil)}_{V}$,
where $\lceil x\rceil$ is the smallest integer greater than or equal to $x$.
7:The bootstrap percentile $(1-\alpha)$-confidence interval for $\theta_{V}$ is
$\bigl[\hat{\theta}^{*}_{\alpha/2},\,\hat{\theta}^{*}_{1-\alpha/2}\bigr].$ (3.4)
Another possible application of the methodology discussed in Section 2 is the assessment of a diagnostic test based on imaging features. Each patient is represented by a feature vector $\mathbb{x}\in\Re^{q}$, and the diagnostic procedure assigns a real-valued score via a function $s:\Re^{q}\to\Re$. Patients are to be classified into three ordered categories (e.g., healthy, mildly diseased, severely diseased) on the basis of this score.
Let $\mathbb{X}$, $\mathbb{Y}$, and $\mathbb{Z}$ denote independent observations from the healthy, mildly diseased, and severely diseased populations, respectively. The parameter of interest is
$\theta_{s}=P\big(s(\mathbb{X})<s(\mathbb{Y})<s(\mathbb{Z})\big),$
which is the probability that a randomly selected triplet, one patient from each group, is correctly ordered by the scoring function $s(\cdot)$. A natural point estimator of $\theta_{s}$ provides a summary measure of the accuracy of the diagnostic test, and one may also be interested in constructing a confidence interval for $\theta_{s}$.
4. Simulation